# A Tutte-type canonical decomposition of 3- and 4-connected graphs

### v2

@@ 9.3. Settings
The following settings are supported by Figure˜9.4.
@@ 9.4. N-Paths lemmata
By Lemma˜9.14, there is $i\in\{1,2\}$ such that $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(i1)t(i)t(i-1)t(i). Moreover, the torso $H_{t(i)}$ is not a 4-cycle. We consider two cases, supported by Figure˜9.5.
@@ 9.5. Totally-nested tetra-separations from good bags
Case $i<j$. This case is supported by Figure˜9.6. If $H_{t(i)}$ is 3-connected, we let $F_{y}$ by a $y^{W}_{1}$$\pi_{i}$ fan of two paths in $G_{t(i)}$ avoiding the vertex in $\pi_{i-1}-y^{W}_{1}$. Otherwise, by Lemma˜9.27, $H_{t(i)}$ is a triangle avoiding $y^{W}_{2}$ such that both edges incident to $y^{W}_{1}$ in $H_{t(i)}$ exist in $G$, and we let $F_{y}$ denote the $y^{W}_{1}$$\pi_{i}$ fan of two paths defined by these two edges. In both cases, $F_{y}$ avoids $y_{2}$. Similarly, we obtain an $x^{W}_{1}$$\pi_{j-1}$ fan $F_{x}\subseteq G_{t(j)}$ avoiding $x_{2}$. By Lemma˜9.16, we find two disjoint $\pi_{i}$$\pi_{j-1}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j-1)}$. Call these paths $P$ and $Q$. Then $F_{x}\cup P\cup Q\cup F_{y}$ is a union of two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$.
@@
Case $i\geqslant j$. This case is supported by Figure˜9.7. We claim that the corner $BD$ is not potter, and assume for a contradiction that it is. Let $\pi$ denote the unique adhesion-set of the $BD$-Tutte-banana. Since $BD$ is potter, neither corner $BC$ nor $AD$ is potter by ˜4.7. Hence if $y_{1}$ is an edge, it neither dangles from the $C$-link through the $B$-link nor from the $B$-link through the $C$-link, by the (Crossing Lemma). (4.1). So $y^{W}_{1}\in C\smallsetminus D$, and hence $y^{W}_{1}\notin\pi$. Similarly, $x^{W}_{1}\notin\pi$. But this contradicts $i\geqslant j$, so the corner $BD$ is not potter.
@@ 9.6. Splitting stars and torsos
Let $P$ be a path, and let $u,v$ be vertices on $P$. Recall that $uP\mathring{v}$ denotes the subpath of $P$ that starts at $u$ and ends just before the vertex $v$. The statement of the following lemma is supported by Figure˜9.8.

### v3

@@ 9.3. Settings
The following settings are supported by fig.˜9.4.
@@ 9.4. N-Paths lemmata
By lemma˜9.14, there is $i\in\{1,2\}$ such that $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(i1)t(i)t(i-1)t(i). Moreover, the torso $H_{t(i)}$ is not a 4-cycle. We consider two cases, supported by fig.˜9.5.
@@ 9.5. Totally-nested tetra-separations from good bags
Case $i<j$. This case is supported by fig.˜9.6. If $H_{t(i)}$ is 3-connected, we let $F_{y}$ by a $y^{W}_{1}$$\pi_{i}$ fan of two paths in $G_{t(i)}$ avoiding the vertex in $\pi_{i-1}-y^{W}_{1}$. Otherwise, by lemma˜9.27, $H_{t(i)}$ is a triangle avoiding $y^{W}_{2}$ such that both edges incident to $y^{W}_{1}$ in $H_{t(i)}$ exist in $G$, and we let $F_{y}$ denote the $y^{W}_{1}$$\pi_{i}$ fan of two paths defined by these two edges. In both cases, $F_{y}$ avoids $y_{2}$. Similarly, we obtain an $x^{W}_{1}$$\pi_{j-1}$ fan $F_{x}\subseteq G_{t(j)}$ avoiding $x_{2}$. By lemma˜9.16, we find two disjoint $\pi_{i}$$\pi_{j-1}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j-1)}$. Call these paths $P$ and $Q$. Then $F_{x}\cup P\cup Q\cup F_{y}$ is a union of two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$.
@@
Case $i\geqslant j$. This case is supported by fig.˜9.7. We claim that the corner $BD$ is not potter, and assume for a contradiction that it is. Let $\pi$ denote the unique adhesion-set of the $BD$-Tutte-banana. Since $BD$ is potter, neither corner $BC$ nor $AD$ is potter by ˜4.7. Hence if $y_{1}$ is an edge, it neither dangles from the $C$-link through the $B$-link nor from the $B$-link through the $C$-link, by the (Crossing Lemma). (4.1). So $y^{W}_{1}\in C\smallsetminus D$, and hence $y^{W}_{1}\notin\pi$. Similarly, $x^{W}_{1}\notin\pi$. But this contradicts $i\geqslant j$, so the corner $BD$ is not potter.
@@ 9.6. Splitting stars and torsos
Let $P$ be a path, and let $u,v$ be vertices on $P$. Recall that $uP\mathring{v}$ denotes the subpath of $P$ that starts at $u$ and ends just before the vertex $v$. The statement of the following lemma is supported by fig.˜9.8.